A school principal claims that more than 67% of the 700 students of their school prefer printed modules. Test the hypothesis using 95% level of confidence.


1
Expert's answer
2022-06-03T09:13:14-0400

"H_0:p\le0.67"

"H_a:p>0.67"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Evidence:

Based on the information provided, the significance level is "\alpha = 0.05

," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z >1.6449\}."

The z-statistic is computed as follows:



"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{X/700-0.67}{\sqrt{\dfrac{0.67(1-0.67)}{700}}}"